11. A car has two wipers which do not overlap. Each wiper has a blade of length 25 cm sweeping through an angle of 115°. Find the total area cleaned at each sweep of the blades.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given, Radius (r) = 25 cm Sector angle (θ) = 115° Since there are 2 blades, The total area of the sector made by wiper = 2×(θ/360°)×π r2 = 2×(115/360)×(22/7)×252 = 2×158125/252 cm2 = 158125/126 = 1254.96 cm2